OpenStudy (anonymous):

5/2=x/6 what is the answer to this solution

jimthompson5910 (jim_thompson5910):

5/2=x/6 5*6=2x ... Cross multiply Can you take it from here?

OpenStudy (anonymous):

yes

jimthompson5910 (jim_thompson5910):

Alright great. Tell me what you get.

OpenStudy (anonymous):

15

jimthompson5910 (jim_thompson5910):

Good you got it